# Machine Learning Message-Passing for the Scalable Decoding of QLDPC Codes

### v1

@@ Abstract
We present Astra, a novel and scalable decoder using graph neural networks. Our decoder works similarly to solving a Sudoku puzzle of constraints represented by the Tanner graph. In general, QLDPC decoding is based on belief propagation (BP, a variant of message-passing) and requires time intensive post-processing methods such as ordered statistics decoding (OSD). Without using any post-processing, Astra achieves higher thresholds and better logical error rates when compared to BP+OSD, both for surface codes trained up to distance 11 and Bivariate Bicycle (BB) codes trained up to distance 18. Moreover, we can successfully extrapolate the decoding functionality: we decode high distances by using decoders trained on lower distances. Astra+OSD is faster than BP+OSD. We show that with decreasing physical error rates, Astra+OSD makes progressively fewer calls to OSD when compared to BP+OSD, even in the context of extrapolated decoding. Astra(+OSD) achieves orders of magnitude lower logical error rates for BB codes compared to BP(+OSD). The source code is open-sourced at https://github.com/arshpreetmaan/astra.
@@ I Introduction
Successfully decoding QECC codes plays a significant role in achieving large scale, practical quantum computations. The accuracy (i.e. reduce the probability of logical error) and scalability (i.e. increasing the QECC code distance should not significantly slow down the decoding process) are some of the most important characteristics of a decoder.
There are various QECC decoding algorithms, each having its own trade-off between speed and accuracy. For surface codes, Minimum Weight Perfect Matching (MWPM) has a threshold higher than Union-Find (UF), but is slower than the latter [1, 2, 3]. Decoders are also usually specific to classes of codes [2]: MWPM is usually applied to topological codes, whereas BP+OSD is very often used for QLDPC codes. A unification of graph based decoding algorithms is in [4].
In general, BP does not have a threshold for the surface code and fails to decode most of the time [2, 5]. Thus, a second stage decoder (e.g. OSD) does most of the decoding, and this fact increases the decoding runtime. Finding a (machine learning) method that can successfully decode QLPDC codes in a single stage has been an open problem.
To bridge this gap, we propose a novel machine learning decoder called Astra [6]. Our decoder is based on graph neural networks (GNN) which learn a message-passing algorithm for decoding. Astra can be used for arbitrary code distances, and outperforms BP+OSD (Fig. 1) and BP (Fig. 2). Astra has the following features: 1) works on any family of codes that can be represented by Tanner graphs; 2) is capable of transfer learning; 3) has faster runtimes compared to BP+OSD without sacrificing its accuracy and scalability.
In the following, we discuss how Astra improves the state-of-the-art of BP and machine learning decoders (Section I.1). In Section II we discuss how our decoder is similar to solving Sudoku (Section II.1), we describe the loss function used for training (Section II.2) and the training procedure itself (Section II.4). Section III showcases the performance of our decoder on surface and BB codes, and we conclude by Section IV.
@@ I.1 Related Work
Recent work focused on improving the cost of the second stage decoder (e.g. Ambiguity Clustering  [7] and BP+LSD (Localised Statistics Decoding)  [8]) In contrast, Astra focuses on improving of the first stage of decoding, such that the second stage can be completely avoided.
Independently of our work, the recent work of [9] uses a similar GNN approach to decoding QLDPC codes. However, Astra is capable of extrapolating both decoding and training (Section II), uses an improved loss function to overcome the challenges of learning in the presence of degenerate errors (Section II.2). Finally, Astra is more scalable and faster than BP+OSD and can be trained on commodity GPUs.
@@ I.2 Background
Under the circuit level depolarising noise model the surface code has a $1\%$ threshold  [38, 39, 40] – one of the highest known so far. However, the surface code has a low encoding rate, because it encodes a single logical qubit. To overcome this limitation, high rate quantum LDPC codes were proposed. One of the recent LDPC codes are Bivariate Bicycle (BB) [41]. The BB code encoding 12 logical qubits requires only 288 physical qubits instead of 3000 physical qubits to give same level of error suppression as the surface code. Under circuit level noise, the BB codes have a threshold of $0.8\%$.
@@ I.2.1 Belief Propagation and Message-Passing
Belief propagation [42] (BP) is a general message-passing algorithm that has been extensively applied in the context of iterative decoding of classical Low-Density Parity-Check (LDPC) codes. More recently, BP has been adapted to the setting of quantum error-correcting code (QECC) decoding [43].
@@
BP-based decoding algorithms are efficient in terms of computation time and memory requirements. However, BP is not always guaranteed to converge (see Fig. 2 where pure BP does not have a threshold) to the correct decoding solution. BP can suffer from issues like loopy behavior and local minima  [44, 5, 45, 46].
@@ II.1 Solving Decoding Constraints using GNNs
In a nutshell, instead of decoding using a monolithic, large NN which predicts errors given the syndromes, Astra is converting the decoding problem into the one of learning three local functions f, g and r. The functions work according to the local constraints given by the Tanner graph.
@@ II.2 Loss Function
The large degeneracy (i.e. many errors have the same syndrome measurement) of the QECCs is a serious challenge when training machine learning models. The original Sudoku approach presented by [49] does not include degenerate solutions: there is a unique Sudoku solution if there are at least 17 filled cells in the $9\times 9$ board.
@@
Figure 4: Decoding surface codes with Astra+OSD vs BP+OSD by using OSD0 in the second stage. a) Astra+OSD achieves orders of magnitude better LER than BP+OSD and requires fewer OSD calls; b) Speedups of Astra+OSD vs BP+OSD are obtained because Astra converges more often than BP and, consequently, the OSD stage is called significantly fewer times. This holds even when performing extrapolated decoding with the d11 decoder e.g. for distance 25, at 0.06 error rate, Astra+OSD is 250x faster than BP+OSD.
Table 1: Astra hyperparameter values achieving lowest LER
Code d N Itrs Node f. Edge f. LR WD msgnet size msg dropout gru dropout Training set Test set
Surface 3 30 500 500 1e-4 1e-4 512 0.05 0.05 $0.1510^{5}@$ 0.15 $0.0510^{4}@$ 0.05
- 5 30 500 500 1e-4 1e-4 512 0.05 0.05 $0.152*10^{5}@$ 0.15 $0.0510^{4}@$ 0.05
- 7 45 500 500 1e-4 1e-4 512 0.05 0.15 $0.154*10^{5}@$ 0.15 $0.0510^{4}@$ 0.05
- 9 100 500 500 1e-4 1e-4 512 0.05 0.05 $0.155*10^{5}@$ 0.15 $0.0510^{4}@$ 0.05
- 11 100 500 500 1e-4 1e-4 512 0.05 0.05 $0.155*10^{5}@$ 0.15 $0.0510^{4}@$ 0.05
BB 6 30 50 50 1e-4 1e-4 128 0.05 0.05 $0.152*10^{4}@$ 0.15 $0.0510^{3}@$ 0.05
- 12 30 50 50 1e-4 1e-4 128 0.05 0.05 $0.182*10^{4}@$ 0.18 $0.0810^{3}@$ 0.08
- 18 45 50 50 1e-4 1e-4 128 0.05 0.05 $0.1510^{5}@$ 0.15 $0.15*10^{3}@$ 0.1
@@ II.3 Correlated Decoding
We will benchmark Astra using the code capacity depolarising noise model (Section I.2), and there are three possible Pauli errors:X, Z and Y (decomposed into X and Z errors). Most of the decoders do independent (un-correlated) decoding for X and Z channels, but this technique fails to decode correlated Y errors and results in sub-optimal decoding performance.
To learn the message-passing decoding of correlated errors, we perform 1-bit hot encoding for errors where X errors are represented by 1, Z errors are 2 and Y errors are 3. This helps the decoder to learn the relationship between different syndrome qubit flips generated by different types of errors.
@@ II.4 Training
Astra is currently implemented and trained using PyTorch. During training, we use disjoint training and test datasets composed of error-syndrome tuples. We performed hyperparameter search using optuna [52]. The hyperparameters of the successful trainings are tabulated in the Appendix in Table 1. The training is stopped when: 1) the LER of the decoder is substantially lower than the physical error rate for the test set, and 2) the loss over the test set converges to zero.
@@
We also did extrapolated training of large distances for surface code and BB code. We trained d11 surface code starting the training from the d9 trained model. Similarly, we trained BB d12 decoders starting from d6, and BB d18 starting from d12. We observed massive speed-ups of the training procedure without sacrificing the LER of the decoders.
@@ III.2 Decoding Bivariate Bicycle Codes
IV Conclusion
We introduced Astra, a novel, scalable graph neural network decoder that is trained on the Tanner graph of QLDPC codes, such as the surface and BB codes. Our decoder does not require post-processing in order to outperform BP+OSD. When using OSD, Astra+OSD is significantly faster than BP+OSD. Our decoder can be trained easily on commodity GPUs and exhibits features of transfer learning. We are using Astra to decode distances which are higher than the ones used during training, and we are still outperforming BP+OSD. Future work will focsu on extending the current results to circuit-level noise.

### v2

@@ Abstract
We present Astra, a novel and scalable decoder using graph neural networks. Our decoder works similarly to solving a Sudoku puzzle of constraints represented by the Tanner graph. In general, Quantum Low Density Parity Check (QLDPC) decoding is based on Belief Propagation (BP, a variant of message-passing) and requires time intensive post-processing methods such as Ordered Statistics Decoding (OSD). Without using any post-processing, Astra achieves higher thresholds and better logical error rates when compared to BP+OSD, both for surface codes trained up to distance 11 and Bivariate Bicycle (BB) codes trained up to distance 18. Moreover, we can successfully extrapolate the decoding functionality: we decode high distances (surface code up to distance 25 and BB code up to distance 34) by using decoders trained on lower distances. Astra+OSD is faster than BP+OSD. We show that with decreasing physical error rates, Astra+OSD makes progressively fewer calls to OSD when compared to BP+OSD, even in the context of extrapolated decoding. Astra(+OSD) achieves orders of magnitude lower logical error rates for BB codes compared to BP(+OSD). The source code is open-sourced at https://github.com/arshpreetmaan/astra.
@@ I Introduction
The successful decoding of Quantum Error Correction (QEC) codes plays a significant role in achieving large scale, practical quantum computations. The accuracy (i.e. reduce the probability of logical error) and scalability (i.e. increasing the QEC code distance should not significantly slow down the decoding process) are some of the most important characteristics of a decoder.
There are various QEC decoding algorithms, each having its own trade-off between speed and accuracy. For surface codes, Minimum Weight Perfect Matching (MWPM) [1, 2] has a threshold higher than Union-Find (UF)  [3], but is slower than the latter [4, 5, 6]. Decoders are also usually specific to classes of codes [5]: MWPM is usually applied to topological codes, whereas Belief Propagation + Ordered Statistics Decoding (BP+OSD)  [7, 8] is very often used for QLDPC codes. A unification of graph based decoding algorithms is in [9].
In general, BP does not have a threshold for the surface code and fails to decode most of the time [5, 8]. Thus, a second stage decoder (e.g. OSD) [7] does most of the decoding, and this fact increases the decoding runtime. Finding a (machine learning) method that can successfully decode QLPDC codes in a single stage has been an open problem.
To bridge this gap, we propose a novel machine learning decoder called Astra [10]. Our decoder is based on graph neural networks (GNN) which learn a message-passing algorithm for decoding. Astra can be used for arbitrary code distances, and outperforms BP+OSD (Fig. 1) and BP (Fig. 2). Astra has the following features: 1) works on any family of codes that can be represented by Tanner graphs; 2) is capable of transfer learning; 3) has faster runtimes compared to BP+OSD while improving on accuracy.
In the following, we discuss how Astra improves the state-of-the-art of BP and machine learning decoders (Section I.1). In Section II we discuss how our decoder is similar to solving Sudoku (Section II.1), we describe the loss function used for training (Section II.2) and the training procedure of Astra (Section II.4). Section III illustrates the performance of our decoder on surface and BB codes. We conclude and formulate future work in Section V.
@@ I.1 Related Work
Recent work focused on improving the cost of the second stage decoder e.g. BP+CB (closed branch) [11], Ambiguity Clustering  [12] and BP+LSD (Localised Statistics Decoding)  [13]). In contrast, Astra focuses on improving the first stage of decoding, such that the second stage can be completely avoided.
Independent of our work, the recent work of [14] uses a similar GNN approach to decoding QLDPC codes. However, in contrast, Astra is capable of extrapolating both decoding and training (Section II), uses an improved loss function to overcome the challenges of learning in the presence of degenerate errors (Section II.2). Finally, Astra is more scalable and faster than BP+OSD and can be trained on commodity GPUs.
@@ I.2 Background
Under the circuit level depolarising noise model the surface code has a $1\%$ threshold  [43, 44, 45] – one of the highest known so far. However, the surface code has a low encoding rate, because it encodes a single logical qubit. To overcome this limitation, high rate quantum LDPC codes were proposed. One of the recent LDPC codes are Bivariate Bicycle (BB) [46]. The BB code encoding 12 logical qubits requires only 288 physical qubits instead of 3000 physical qubits to give same level of error suppression as the surface code. The BB codes, under circuit level noise, have a high pseudo-threshold of $0.8\%$ point when logical error rate goes below physical error rate.
@@ I.2.1 Belief Propagation and Message-Passing
Belief propagation [47] (BP) is a specific message-passing algorithm that has been extensively applied in the context of iterative decoding of classical Low-Density Parity-Check (LDPC) codes. More recently, BP has been adapted to the setting of QECC decoding (e.g. [48]).
@@
BP-based decoding algorithms are efficient in terms of computation time and memory requirements. However, BP is not always guaranteed to converge (see Fig. 2 where pure BP does not have a threshold) to the correct decoding solution. BP can suffer from issues like loopy behavior and local minima  [49, 8, 50, 51, 7].
@@ II.1 Solving Decoding Constraints using GNNs
In a nutshell, instead of decoding using a monolithic, large NN which predicts errors given the syndromes, Astra is converting the decoding problem into the one of learning three local functions f, g and r. The functions work according to the local constraints from the Tanner graph.
@@ II.2 Loss Function
The large degeneracy (i.e. many errors generate the same syndrome) of the QECCs is a serious challenge when training machine learning models. The original Sudoku approach presented by [54] does not include degenerate solutions: there is a unique Sudoku solution if there are at least 17 filled cells in the $9\times 9$ board.
@@
Figure 4: Decoding surface codes with Astra+OSD vs BP+OSD by using OSD0 in the second stage. a) Astra+OSD achieves orders of magnitude better LER than BP+OSD and requires fewer OSD calls; b) Speedups of Astra+OSD vs BP+OSD are obtained because Astra converges more often than BP and, consequently, the OSD stage is called significantly fewer times. This holds even when performing extrapolated decoding with the d11 decoder e.g. for distance 25, at 0.06 error rate, Astra+OSD is 400x faster than BP+OSD.
Table 1: Astra hyperparameter values achieving lowest LER.
Code d Num. of Iterations Node features Edge features Msgnet size Training set Test set
Surface 3 30 500 500 512 $0.1510^{5}@$ 0.15 $0.0510^{4}@$ 0.05
- 5 30 500 500 512 $0.152*10^{5}@$ 0.15 $0.0510^{4}@$ 0.05
- 7 45 500 500 512 $0.154*10^{5}@$ 0.15 $0.0510^{4}@$ 0.05
- 9 100 500 500 512 $0.155*10^{5}@$ 0.15 $0.0510^{4}@$ 0.05
- 11 100 500 500 512 $0.155*10^{5}@$ 0.15 $0.0510^{4}@$ 0.05
BB 6 30 50 50 128 $0.152*10^{4}@$ 0.15 $0.0510^{3}@$ 0.05
- 12 30 50 50 128 $0.182*10^{4}@$ 0.18 $0.0810^{3}@$ 0.08
- 18 45 50 50 128 $0.1510^{5}@$ 0.15 $0.15*10^{3}@$ 0.1
@@ II.3 Correlated Decoding
We benchmark Astra using the code capacity depolarising noise model (Section I.2), and there are three possible Pauli errors: X, Z and Y (decomposed into X and Z errors). Most of the decoders do independent (un-correlated) decoding for X and Z channels, but this technique fails to decode correlated Y errors and results in sub-optimal decoding performance.
To learn the message-passing decoding of correlated errors, we use 1-bit hot encoding for errors where X errors are represented by 1, Z errors are 2 and Y errors are 3. This helps the decoder to learn the relationship between different syndrome qubit flips generated by different types of errors.
@@ II.4 Training
Astra is currently implemented and trained using PyTorch. During training, we use disjoint training and test datasets composed of error-syndrome tuples. We performed hyperparameter search using optuna [57]. The hyperparameters of the successful trainings are tabulated in Table 1. Additionally, we used a learning rate of $41e-$ 4, a dropout rate of $0.05$ in the message net layer. The training is stopped when: 1) the LER of the decoder is substantially lower than the physical error rate for the test set, and 2) the loss over the test set converges to zero.
@@
We also executed extrapolated training of large distances for surface code and BB code. We trained d11 surface code starting the training from the d9 trained model. Similarly, we trained BB d12 decoders starting from d6, and BB d18 starting from d12. We observed massive speed-ups of the training procedure without sacrificing the LER of the decoders.
@@ III.2 Decoding Bivariate Bicycle Codes
IV Discussion
For all the simulation data with the BP decoder we used the serial scheduling, a maximum of 100 iterations, the minimum sum algorithm and a scaling factor of 0.75 for BB codes and a scaling factor of 1.0 for surface codes. For Astra we used the flooding scheduling i.e. sending all the messages at once. In general, BP has better decoding performance when using serial scheduling rather than flooding. One can, in principle, implement different scheduling methods with GNN too, be it serial, parallel or random. We did not experiment with these but we expect that changing the scheduling might help the Astra decoder even better like in the case of BP.
V Conclusion
We introduced Astra, a novel, scalable graph neural network decoder that is trained on the Tanner graph of QLDPC codes, such as the surface and BB codes. Our decoder does not require post-processing in order to outperform BP+OSD. When using OSD, Astra+OSD is significantly faster than BP+OSD. Our decoder can be trained easily on commodity GPUs and exhibits features of transfer learning. We are using Astra to decode distances which are higher than the ones used during training, and we are still outperforming BP+OSD. Future work will focus on extending the current results to circuit-level noise.